# Impact of domain reduction techniques in polynomial optimization: A computational study

### v1

@@ 5 Conclusions
We have studied the impact of several approaches to enhance the efficacy of domain reduction functionalities within the specific context of the RLT technique for polynomial optimization problems. The presented results showed that, although all these enhancements lead to performance improvements, the overall impact may sometimes be relatively small. Interestingly, the largest impact comes from the approaches to select the branching point which are, by far, the simplest and easiest to implement among all the enhancements under consideration. We believe these kind of insights on the trade-offs between performance improvement and implementation costs are very relevant for the design of more efficient optimization solvers, since they can guide the implementation efforts of the developing teams of state-of-the-art solvers towards the most promising directions. When a new technique or algorithmic enhancement is first studied, it is natural to analyze is impact in isolation, in order to better quantify its impact. Yet, because of this isolated analyses, it is often hard to assess which of a series of potential enhancements should be prioritized when designing a new solver or when evolving an already existing one, since the impacts individually reported for each of them might no be easy to compare to one another.
Following recent trends in the field, we also built upon the learning framework developed in [Ghaddar:2023] to show that machine/statistical learning can be used to further improve performance. The promising results obtained in Section 4.3, with the direct application of the methodology in [Ghaddar:2023], suggests that one could get even better results with tailor-made modifications of it. In particular, a relatively simple avenue for future research would be to extend the set of features used for the learning, including additional ones that may be suitable for domain reduction.
@@ Acknowledgments
Grant PID2021-124030NB-C32 funded by MICIU/AEI/10.13039/501100011033/ and by ERDF/EU. This research was also funded by Grupos de Referencia Competitiva ED431C-2021/24 from the Consellería de Cultura, Educación e Universidades, Xunta de Galicia. Ignacio Gómez-Casares acknowledges the support from the Spanish Ministry of Education through FPU grant 20/01555.
@@ Appendix A Additional results for FBBT enhancements
Table 9 presents a more comprehensive summary of the numerical results obtained when studying different variations of the FBBT approaches described in Section 3.2. We study configurations in which the two approaches are applied at all nodes or just at the root node and at nodes of pre-specified depths. Note that, under the approach FBBT, the cuts from FBBT are never applied by themselves, since multiples of 50 are also multiples of 10. In order to assess whether or not this might impact performance, we also studied FBBT, under which both families of cuts are jointly applied at the same node much less often.
@@
FBBT | 281 | |||||
FBBT | 286 | |||||
FBBT | 287 | |||||
FBBT | 287 | |||||
FBBT | 287 | |||||
FBBT | 287 | |||||
FBBT | 284 | |||||
FBBT | 284 | |||||
FBBT | 284 | |||||
FBBT | 286 | |||||
FBBT | 286 | |||||
FBBT | 287 |

### v2

@@ 5 Conclusions
We have studied the impact of various approaches to improve the effectiveness of domain reduction functionalities in the context of the RLT technique for solving polynomial optimization problems. Although all of these enhancements lead to performance improvements, the results showed that the overall impact may sometimes be relatively small. Interestingly, the approaches to select the branching point have the largest impact and they are, by far, the simplest and easiest to implement among all the enhancements under consideration. We believe that insights into the trade-offs between performance improvement and implementation costs are relevant to the design of more efficient optimization solvers, because they can guide the implementation efforts of state-of-the-art solver development teams toward the most promising directions. When a new technique or algorithmic enhancement is first studied, it is natural to analyze its impact in isolation to better quantify it. However, precisely because of this isolation, it is often difficult to determine which potential enhancements should be prioritized when designing or evolving a solver, as the reported impacts may not be easily comparable.
Following recent trends in the field, we built upon the learning framework developed in [Ghaddar:2023] to demonstrate the potential of machine/statistical learning to enhance performance further. The promising results obtained in Section 4.3, with the direct application of the methodology in [Ghaddar:2023], suggest that one could achieve even better results with modifications tailored to specific applications. A relatively simple avenue for future research would be to extend the set of features used for learning to include additional features suitable for domain reduction.
@@ Acknowledgments
Project PID2021-124030NB-C32 funded by MICIU/AEI/10.13039/501100011033/ and by ERDF/EU. This research was also funded by Grupos de Referencia Competitiva ED431C-2021/24 from the Consellería de Cultura, Educación e Universidades, Xunta de Galicia. Ignacio Gómez-Casares acknowledges the support from the Spanish Ministry of Education through FPU grant 20/01555.
Conflict of interest
The authors declare no competing interests.
@@ Appendix A Additional results for FBBT enhancements
Table 10 presents a more comprehensive summary of the numerical results obtained when studying different variations of the FBBT approaches described in Section 3.2. We study configurations in which the two approaches are applied at all nodes or just at the root node and at nodes of pre-specified depths. Note that, under the approach FBBTOB,50+NB,10, the cuts from FBBTOB are never applied by themselves, since multiples of 50 are also multiples of 10. In order to assess whether or not this might impact performance, we also studied FBBTOB,50+NB,9, under which both families of cuts are jointly applied at the same node much less often.
When looking at the performance of the approaches in which FBTOB or FBTNB are applied at all nodes, we can see that the reductions on the sizes of the resulting trees do not compensate the computational overhead. In particular, when comparing FBBTOB,all with FBBTOB,25 and FBBTOB,50, the value of Nodes shows that FBBTOB,all yields slightly smaller trees than the other approaches but, on the other hand, Time becomes much larger. A similar observation applies to FBBTNB,all with respect to FBBTNB,5 and FBBTNB,10, although the increase in Time is less pronounced.
@@
| FBBTOB,all | 281 | |||||
| FBBTOB,25 | 286 | |||||
| FBBTOB,50 | 287 | |||||
| FBBTNB,all | 287 | |||||
| FBBTNB,5 | 287 | |||||
| FBBTNB,10 | 287 | |||||
| FBBTOB,25+NB,5 | 284 | |||||
| FBBTOB,25+NB,9 | 284 | |||||
| FBBTOB,25+NB,10 | 284 | |||||
| FBBTOB,50+NB,5 | 286 | |||||
| FBBTOB,50+NB,9 | 286 | |||||
| FBBTOB,50+NB,10 | 287 |